### v1

@@ 1 Introduction
Reasoning tasks in chemistry are fundamental yet notoriously challenging, requiring models to integrate multiple layers of chemical knowledge and logical deduction [1, 2]. While traditional chemoinformatics approaches rely primarily on supervised architectures optimized for specific tasks, they lack generalization and human-like reasoning capacities, instead often performing as highly specialized pattern recognition systems [3, 4]. Recently, reinforcement learning (RL) driven frameworks [5, 6, 7, 8] have shown promising advances in generating sophisticated emergent reasoning capabilities without explicit step-level supervision, achieving remarkable results across general-purpose domains like math and coding. Nevertheless, independent follow-ups have shown that such capabilities do not simply appear, but emerge instead as amplified patterns already existing in the base model’s output distribution —even if with low likelihoods [9, 10]. Consequently, whether RL succeeds on a new domain depends crucially on the latent solvability of the tasks for that specific base model.
Chemistry presents a severe stress test for this premise. Unlike arithmetic or programming, chemical problems combine highly specialized symbol systems [11] (e.g. Simplified Molecular Input Line Entry System –SMILES, IUPAC names) with domain-specific physical constraints (valence, stereochemistry). Off-the-shelf LLMs typically fail to generate syntactically valid SMILES, let alone perform any tasks involving SMILES manipulation and generation [12]. Empirically, we find that direct application of RL methods to such models fails: the reward signal vanishes because the correct answer never appears in the candidate set, except for the simpler examples.
@@ 2.1 Post-training methods for reasoning
The standard recipe for aligning LLMs augments supervised fine-tuning (SFT) with reinforcement learning from human or synthetic feedback (RLHF/RLAIF) [13]. While RLHF reliably improves helpfulness and stylistic alignment, it is often insufficient for multi-step reasoning. Subsequent work therefore introduced chain-of-thought distillation [14, 15], and tree search with self-consistency [16]. A recent, influential result by Guo et al. [5] showed that rule-based rewards can unlock strong mathematical and coding skills, provided that the base model already allocates non-negligible probability mass to correct answers. Independent analyses confirmed that RL mainly acts as an amplifier: it can only surface solutions that lie somewhere in the base distribution [17]. For weaker bases, SFT on traces generated by a larger model often outperforms RL [5]. Our work adopts this "RL as amplifier" view and asks what pre-training conditions make chemical problems latently solvable so that RL can succeed.
@@ 2.2 Chemical language modeling
Language models have been adapted and used for a range of chemical tasks [18]. These models typically operate on linearized molecular strings, such as SMILES [11], SELFIES [19], or IUPAC names. Masked pre-training approaches (ChemBERTa [20], MolBERT [21]) learn molecular fingerprints that are useful for QSAR, whereas the Molecular Transformer family targets forward and retrosynthesis prediction [3, 22, 23]. More recently, LLMs have been adapted and applied for tackling chemical tasks [24, 25, 26, 27]; extending general LLMs for molecule generation, property prediction, and Q&A. Other works have adapted LLMs for use as chemistry agents, integrating robotic labs and other tools [28, 29], hypothesis generation [30, 31, 32], and more recently, workflows have been designed for molecular design and synthesis planning [33, 12, 34, 35].
@@ 2.3 Mid-stage domain adaptation
Continuing pre-training on an in-domain corpus—often called domain-adaptive pre-training (DAPT) or continued pre-training (CPT)—has become the dominant recipe for turning a general LLM into a domain specialist. Early successes such as BioMegatron for biomedicine [36], Meditron [37], Legal-BERT [38], and Code-Llama for programming [39] demonstrated sizable gains with only a few billion extra tokens. A recent wave of work scales the idea to scientific domains: (1) AdaptLLM [40] shows that a 7B parameters model, after just 10–15B of financial tokens, rivals BloombergGPT (50B parameters) on in-domain QA; (2) Tag-LLM [41] and Efficient-CPT [42] report similar jumps while using parameter-efficient adapters; (3) SciLitLLM [43] uses a 12.7B token corpus of textbooks and full-text papers and beats much larger baselines on scientific-literature understanding; (4) domain-specific studies in materials science [44], radiation oncology [45], Japanese finance [46], and cybersecurity [47] confirm that CPT injects latent domain knowledge that survives further instruction tuning. However, two caveats emerge: CPT can erode zero-shot prompting ability if done naively [40], and very small models (<< 2B parameters) often fail to develop new capabilities even after extensive CPT [44, 48]. Crucially, none of these works evaluate whether the adapted model becomes latently solvable for multi-step reasoning tasks that reinforcement learning could later amplify.
The chemical domain remains comparatively under-explored. ChemBERTa-2 [49] continues a BERT-style encoder on $\sim$1B SMILES tokens and improves fingerprint-style QSAR, while ChemLLM [25] , DARWIN [50], and SciDFM [51], LlaSMol [52] incorporate reaction patents or literature but still operate in a single-shot, pattern-recognition regime.
@@ 2.4 LLM capability diagnostics
Benchmark accuracy and perplexity offer only coarse snapshots of a model; they ignore the richer signal contained in the full conditional probability distribution. Holistic evaluation suites such as HELM [53] and LiveBench [54] log likelihoods but still aggregate them into single numbers. Probability-based intrinsic probes provide finer insight. BLiMP minimal pairs measure grammatical preference gaps [55], an idea later reused to analyze in-context learning brittleness [56] and out-of-domain (OOD) intent detection [57]. For factual QA, calibration studies show that token-level probabilities reveal when models "know what they know" [58, 59]. Distributional uncertainty metrics now underpin OOD detection [60], self-correction pipelines [61], and medical-reasoning assessment [62]. Pezeshkpour [63] and Wang et al. [64] formalize diagnostics as distribution-matching problems using KL divergence or Wasserstein distance, while [65] links dispersion measures to downstream robustness.
@@ 3.1 Prerequisite 1: Symbolic Competence
To assess the symbolic competence of models, we compute the likelihood of generating a given sequence, in our case, a set of SMILES strings. We use a dataset of 10,000 molecules obtained from PubChem [66], and use the following definitions to compute a symbolic competence score.
@@ 3.2 Prerequisite 2: Latent Chemical Knowledge
With this in mind, we aim to assess the latent chemical knowledge of a given base model. As a proxy to this, we adopt the same strategy as that we use with the symbolic data, by measuring the Chemical-Competence Score (CCS), defined as the difference in the distributions of mean ranks between factually correct chemical statements and wrong ones. Given a list of chemical statements, such as the SMolInstruct Molecule Description subset [52] and descriptions generated by Robocrystallographer [67], we generate corrupted data by randomly swapping one sentence from each original statement with that from another randomly chosen statement in the pool.
@@ 3.3.1 Supervised fine-tuning on reasoning traces
[68] showed that small base models can be trained with SFT on reasoning traces, resulting in small reasoning models that mimic the behavior demonstrated in the SFT training data, even if such data does not directly target the specific downstream task the models are evaluated on. The reason is that SFT transfers the response style and not only the task-specific capabilities, thus serving as an amplifier of latent knowledge. Following this, some reasoning traces were distilled from DeepSeek-R1 and used to perform SFT on our pretrained models. We generated $\sim$ 600,000 solutions for two canonical tasks: IUPAC $\to$ SMILES and SMILES $\to$ IUPAC, based on PubChem compounds.
@@ 3.3.2 Reinforcement learning from Scientific Feedback
Following recent works [69], we adopt Reinforcement Learning from scientific feedback (RLSF) as a post-training method for our models. In this context, models are trained online with rule-based scientific rewards that depend entirely on the final outcome. The goal of this type of training, as exemplified in previous works [69] is to encourage the model to achieve good results on the training tasks, while developing intermediate strategies to achieve this, that might involve reasoning.
@@ 4.1 Datasets
To construct the pretraining data, we used the data mixture as described in Table 1, obtained and processed as described in Appendix D. For molecular data, all text underwent preprocessing to interleave SMILES with text whenever a molecule name appeared (e.g., IUPAC, common name, short form, etc.), similar to [70]. We additionally generated a synthetic dataset using RDKit [71] with extracted molecular properties (such as QED, TPSA, etc.) filled into templates. For inorganic materials, we applied analogous pre-processing to interleave composition formulas and crystallographic information (e.g., space groups, crystal systems) with descriptive text. We also generated synthetic datasets containing computed properties from density functional theory calculations (such as formation energy, band gap, bulk modulus, etc.) using structured templates. Furthermore, we include a "replay" dataset aiming to preserve the model’s natural language abilities while furthering its learning about chemical and materials knowledge. We chose the Qwen2.5-3B base model to perform the pretraining for 3 epochs.
@@
For SFT, we used question-answering (QA) training examples derived from two domains: (1) for organic molecules, we employed SmolInstruct [72], utilizing the SMILES $\leftrightarrow$ IUPAC and molecule captioning subsets; (2) for inorganic crystals, we employed MPtrj [73], utilizing structure to energy/force prediction and initial to relaxed structure mapping subsets, with reasoning traces constructed from intermediate ionic steps to capture the DFT relaxation pathway. Dataset statistics are summarized in Tables 1.Additionally, we incorporated MMLU and Chain-of-Thought (CoT) reasoning traces from DeepSeek-R1, which were preprocessed to maintain coherence with our pretraining data. In this phase, we also expanded the model’s context window from 4,096 to 8,192 tokens to accommodate longer reasoning sequences. The pretrained model underwent SFT for approximately 8 epochs, continuing until the previously observed loss spikes were fully mitigated.
@@ 5 Results
As an initial downstream test of our pipeline’s performance, we use ChemBench [74] to evaluate the general chemistry knowledge of LLMs; the results are shown in Table 2. This evaluation is a first check to ensure our pipeline consistently improves performance on an important public benchmark for chemistry, before we begin studying the downstream effects of MiST on RL training. The results show an overall +10% improvement in performance bringing it up to 45% which, better than GPT-3.5-Turbo and Llama-3.1-8B-Instruct [4], models at least 2x larger than our resulting 3B model. The role of MiST is particularly important in Organic, Inorganic, and General Chemistry, with improvements of up to 6-7% over the Qwen+SFT baseline, and more than 11% over the base (instruction-tuned) model, suggesting benefits of both post-training stages on model’s chemical performance. These results serve already as diagnostic measures of success of a given mid-training methodology, and serve as a basis to select models for the following stage of RLSF, with the goal of enhancing reasoning and problem-solving capabilities.
@@
We then proceed to evaluate model’s capacity of learning in an online setting from verifiable scientific rewards through RLSF. As explained in Section A, we implement a number of chemical tasks that are suitable for reasoning, and for which verifiable scientific rewards can be defined (see Appendix A). Several models were trained on this setting and, in the following, we evaluate task performance as a function of the base model used. We employ two different inference techniques to generate the results reported here, namely using System-1 (direct answer) and System-2 (employing reasoning) thinking, as defined by [75]. We do this by appending the tags "<answer>" or "<reasoning>" respectively upon generation, which induces models into either type of thinking. The results shown in Table 2 show the performances of our models across the multiple tasks defined in Section A, along with the diagnostic metrics defined in Section 3.1.
@@
Previous works have built models for similar tasks [25, 6, 76]. In Table 2, we compare against a baseline of similar size, ChemLLM, a model trained starting from InternLM2-Base-7B [77]. Our results show that, on the tasks evaluated, ChemLLM is outperformed by our RL-tuned 3B models. The advantage is especially striking in the tasks of reaction naming and and reaction prediction. This contrasts with the results presented in the original publication, reporting over 88% 5-shot accuracy on retrosynthesis which could not be reproduced.
@@ 5.2 SCS as a predictive framework
A main claim of this paper is that it is important to develop diagnostic metrics to help prognosticate the performance of models before RL is applied, similar to the role of scaling laws in pre-training [78]. Here we show that pre-RL SCS is effective for this. As shown in Figure 4, pre-RL SCS reliably predicts post-RL success of models. It is observed that, on the tasks evaluated, low pre-RL SCS predictably yields incapable models, while high SCS leads to better models, especially when the models are trained with RL on the tasks being evaluated. We find strong correlations between pre-RL SCS and post-RL performance, namely $\rho=0.64$ for reaction prediction and $\rho=0.60$ for IUPAC translation across both 3B and 7B models. This provides a quantitative empirical framework to support future model development: SCS > 1.5 is a necessary threshold for RL to unlock reasoning.
@@
Qwen-2.5-14B [79] 1.161
Qwen-2.5-32B [79] 1.238
Qwen-2.5-72B [79] 1.094
Llama3-70B [80] 1.481
Gemma-2-27B [81] 2.097
Mistral-24B [82] 2.324
@@
As case-study, we use SCS to retrospectively select the best base model to train ether0, the first chemical reasoning model [6]. We selected a range of open-source base LLMs across a variety of providers and sizes between 14B and 70B, and measured SCS. The results in Table 5 show that Mistral-24B would be the best option, followed by Gemma-2-27B, while the rest of models, even some beyond 70B in size, fall below the 1.5 SCS limit established previously. Within a similar size range, and despite its much better scores at most public benchmarks [82], Qwen-2.5-32B is predicted to be a much worse base model for RL in chemistry tasks. These results match the base model selection from ether0 [6], that effectively used Mistral-24B as a base, further increasing confidence on the predictive power of measures of this type.

### v2

@@ 1 Introduction
Reasoning tasks in chemistry are fundamental yet notoriously challenging, requiring models to integrate multiple layers of chemical knowledge and logical deduction [coley2019graph, alampara2024probing]. While traditional chemoinformatics approaches rely primarily on supervised architectures optimized for specific tasks, they lack generalization and human-like reasoning capacities, instead often performing as highly specialized pattern recognition systems [schwaller2019molecular, mirza2024large]. Recently, reinforcement learning (RL) driven frameworks [guo2025deepseek, ether0, chemdfmr, chemr] have shown promising advances in generating sophisticated emergent reasoning capabilities without explicit step-level supervision, achieving remarkable results across general-purpose domains like math and coding. Nevertheless, independent follow-ups have shown that such capabilities do not simply appear, but emerge instead as amplified patterns already existing in the base model’s output distribution —even if with low likelihoods [guo2023can, flamshepherd2023languagemodelsgeneratemolecules]. Consequently, whether RL succeeds on a new domain depends crucially on the latent solvability of the tasks for that specific base model.
Chemistry presents a severe stress test for this premise. Unlike arithmetic or programming, chemical problems combine highly specialized symbol systems [weininger1988smiles] (e.g. Simplified Molecular Input Line Entry System –SMILES, IUPAC names) with domain-specific physical constraints (valence, stereochemistry). Off-the-shelf LLMs typically fail to generate syntactically valid SMILES, let alone perform any tasks involving SMILES manipulation and generation [bran2025chemical]. Empirically, we find that direct application of RL methods to such models fails: the reward signal vanishes because the correct answer never appears in the candidate set, except for the simpler examples.
@@ 2.1 Post-training methods for reasoning
The standard recipe for aligning LLMs augments supervised fine-tuning (SFT) with reinforcement learning from human or synthetic feedback (RLHF/RLAIF) [lee_rlaif_2024]. While RLHF reliably improves helpfulness and stylistic alignment, it is often insufficient for multi-step reasoning. Subsequent work therefore introduced chain-of-thought distillation [wei2022chain, Li2023SymbolicCD], and tree search with self-consistency [Xie2024MonteCT]. A recent, influential result by guo2025deepseek showed that rule-based rewards can unlock strong mathematical and coding skills, provided that the base model already allocates non-negligible probability mass to correct answers. Independent analyses confirmed that RL mainly acts as an amplifier: it can only surface solutions that lie somewhere in the base distribution [yue2025doesreinforcementlearningreally]. For weaker bases, SFT on traces generated by a larger model often outperforms RL [guo2025deepseek]. Our work adopts this "RL as amplifier" view and asks what pre-training conditions make chemical problems latently solvable so that RL can succeed.
@@ 2.2 Chemical language modeling
Language models have been adapted and used for a range of chemical tasks [caldasramos_review_2025]. These models typically operate on linearized molecular strings, such as SMILES [weininger1988smiles], SELFIES [krenn2020self], or IUPAC names. Masked pre-training approaches (ChemBERTa [chithrananda2020chemberta], MolBERT [Fabian2020MolecularRL]) learn molecular fingerprints that are useful for QSAR, whereas the Molecular Transformer family targets forward and retrosynthesis prediction [schwaller2019molecular, schwaller2020predicting, irwin2022chemformer]. More recently, LLMs have been adapted and applied for tackling chemical tasks [Frey2023NeuralSO, zhang2024chemllmchemicallargelanguage, Jablonka2024LeveragingLL, Xie2023DARWINSD]; extending general LLMs for molecule generation, property prediction, and Q&A. Other works have adapted LLMs for use as chemistry agents, integrating robotic labs and other tools [Bran2023AugmentingLL, boiko2023autonomous], hypothesis generation [moosechem, airesearcher, kosmos], and more recently, workflows have been designed for molecular design and synthesis planning [wang2024molleo, bran2025chemical, synthelite, larcretro].
@@ 2.3 Mid-stage domain adaptation
Continuing pre-training on an in-domain corpus—often called domain-adaptive pre-training (DAPT) or continued pre-training (CPT)—has become the dominant recipe for turning a general LLM into a domain specialist. Early successes such as BioMegatron for biomedicine [shin2020biomegatron], Meditron [meditron], Legal-BERT [chalkidis2020legal], and Code-Llama for programming [roziere2023code] demonstrated sizable gains with only a few billion extra tokens. A recent wave of work scales the idea to scientific domains: (1) AdaptLLM [cheng2023adapting] shows that a 7B parameters model, after just 10–15B of financial tokens, rivals BloombergGPT (50B parameters) on in-domain QA; (2) Tag-LLM [shen2024tag] and Efficient-CPT [xie2024efficientcpt] report similar jumps while using parameter-efficient adapters; (3) SciLitLLM [li2024scilitllm] uses a 12.7B token corpus of textbooks and full-text papers and beats much larger baselines on scientific-literature understanding; (4) domain-specific studies in materials science [lu2025finetuning], radiation oncology [holmes2023radiation], Japanese finance [hirano2024finance], and cybersecurity [bayer2024cysecbert] confirm that CPT injects latent domain knowledge that survives further instruction tuning. However, two caveats emerge: CPT can erode zero-shot prompting ability if done naively [cheng2023adapting], and very small models (<< 2B parameters) often fail to develop new capabilities even after extensive CPT [lu2025finetuning, hsieh2025continual]. Crucially, none of these works evaluate whether the adapted model becomes latently solvable for multi-step reasoning tasks that reinforcement learning could later amplify.
The chemical domain remains comparatively under-explored. ChemBERTa-2 [maziarka2023chemberta2] continues a BERT-style encoder on $\sim$1B SMILES tokens and improves fingerprint-style QSAR, while ChemLLM [zhang2024chemllmchemicallargelanguage] , DARWIN [xie2023darwin], and SciDFM [sun2024scidfm], LlaSMol [yu2024llasmol] incorporate reaction patents or literature but still operate in a single-shot, pattern-recognition regime.
@@ 2.4 LLM capability diagnostics
Benchmark accuracy and perplexity offer only coarse snapshots of a model; they ignore the richer signal contained in the full conditional probability distribution. Holistic evaluation suites such as HELM [liang2022holistic] and LiveBench [white2024livebench] log likelihoods but still aggregate them into single numbers. Probability-based intrinsic probes provide finer insight. BLiMP minimal pairs measure grammatical preference gaps [meister2021beyond], an idea later reused to analyze in-context learning brittleness [zhao2024decision] and out-of-domain (OOD) intent detection [wang2024oodintent]. For factual QA, calibration studies show that token-level probabilities reveal when models "know what they know" [jiang2021know, kadavath2022language]. Distributional uncertainty metrics now underpin OOD detection [liu2023ooddetect], self-correction pipelines [liu2024selfcorrection], and medical-reasoning assessment [li2024mediq]. pezeshkpour2023measuring and wang2024assessing formalize diagnostics as distribution-matching problems using KL divergence or Wasserstein distance, while [ye2024uq] links dispersion measures to downstream robustness.
@@ 3.1 Prerequisite 1: Symbolic Competence
To assess the symbolic competence of models, we compute the likelihood of generating a given sequence, in our case, a set of SMILES strings. We use a dataset of 10,000 molecules obtained from PubChem [pubchem_2025], and use the following definitions to compute a symbolic competence score.
@@ 3.2 Prerequisite 2: Latent Chemical Knowledge
With this in mind, we aim to assess the latent chemical knowledge of a given base model. As a proxy to this, we adopt the same strategy as that we use with the symbolic data, by measuring the Chemical-Competence Score (CCS), defined as the difference in the distributions of mean ranks between factually correct chemical statements and wrong ones. Given a list of chemical statements, such as the SMolInstruct Molecule Description subset [yu2024llasmol] and descriptions generated by Robocrystallographer [ganose2019robocrystallographer], we generate corrupted data by randomly swapping one sentence from each original statement with that from another randomly chosen statement in the pool.
@@ 3.3.1 Supervised fine-tuning on reasoning traces
[deepseek] showed that small base models can be trained with SFT on reasoning traces, resulting in small reasoning models that mimic the behavior demonstrated in the SFT training data, even if such data does not directly target the specific downstream task the models are evaluated on. The reason is that SFT transfers the response style and not only the task-specific capabilities, thus serving as an amplifier of latent knowledge. Following this, some reasoning traces were distilled from DeepSeek-R1 and used to perform SFT on our pretrained models. We generated $\sim$ 600,000 solutions for two canonical tasks: IUPAC $\to$ SMILES and SMILES $\to$ IUPAC, based on PubChem compounds.
@@ 3.3.2 Reinforcement learning from Scientific Feedback
Following recent works [wang_reinforcement_2025], we adopt Reinforcement Learning from scientific feedback (RLSF) as a post-training method for our models. In this context, models are trained online with rule-based scientific rewards that depend entirely on the final outcome. The goal of this type of training, as exemplified in previous works [wang_reinforcement_2025] is to encourage the model to achieve good results on the training tasks, while developing intermediate strategies to achieve this, that might involve reasoning.
@@ 4.1 Datasets
To construct the pretraining data, we used the data mixture as described in Table 1, obtained and processed as described in Appendix D. For molecular data, all text underwent preprocessing to interleave SMILES with text whenever a molecule name appeared (e.g., IUPAC, common name, short form, etc.), similar to [galactica]. We additionally generated a synthetic dataset using RDKit [rdkit] with extracted molecular properties (such as QED, TPSA, etc.) filled into templates. For inorganic materials, we applied analogous pre-processing to interleave composition formulas and crystallographic information (e.g., space groups, crystal systems) with descriptive text. We also generated synthetic datasets containing computed properties from density functional theory calculations (such as formation energy, band gap, bulk modulus, etc.) using structured templates. Furthermore, we include a "replay" dataset aiming to preserve the model’s natural language abilities while furthering its learning about chemical and materials knowledge. We chose the Qwen2.5-3B base model to perform the pretraining for 3 epochs.
@@
For SFT, we used question-answering (QA) training examples derived from two domains: (1) for organic molecules, we employed SmolInstruct [smolinstruct], utilizing the SMILES $\leftrightarrow$ IUPAC and molecule captioning subsets; (2) for inorganic crystals, we employed MPtrj [deng2023chgnet], utilizing structure to energy/force prediction and initial to relaxed structure mapping subsets, with reasoning traces constructed from intermediate ionic steps to capture the DFT relaxation pathway. Dataset statistics are summarized in Tables 1.Additionally, we incorporated MMLU and Chain-of-Thought (CoT) reasoning traces from DeepSeek-R1, which were preprocessed to maintain coherence with our pretraining data. In this phase, we also expanded the model’s context window from 4,096 to 8,192 tokens to accommodate longer reasoning sequences. The pretrained model underwent SFT for approximately 8 epochs, continuing until the previously observed loss spikes were fully mitigated.
@@ 5 Results
As an initial downstream test of our pipeline’s performance, we use ChemBench [ChemBench] to evaluate the general chemistry knowledge of LLMs; the results are shown in Table 2. This evaluation is a first check to ensure our pipeline consistently improves performance on an important public benchmark for chemistry, before we begin studying the downstream effects of MiST on RL training. The results show an overall +10% improvement in performance bringing it up to 45% which, better than GPT-3.5-Turbo and Llama-3.1-8B-Instruct [mirza2024large], models at least 2x larger than our resulting 3B model. The role of MiST is particularly important in Organic, Inorganic, and General Chemistry, with improvements of up to 6-7% over the Qwen+SFT baseline, and more than 11% over the base (instruction-tuned) model, suggesting benefits of both post-training stages on model’s chemical performance. These results serve already as diagnostic measures of success of a given mid-training methodology, and serve as a basis to select models for the following stage of RLSF, with the goal of enhancing reasoning and problem-solving capabilities.
@@
We then proceed to evaluate model’s capacity of learning in an online setting from verifiable scientific rewards through RLSF. As explained in Section A, we implement a number of chemical tasks that are suitable for reasoning, and for which verifiable scientific rewards can be defined (see Appendix A). Several models were trained on this setting and, in the following, we evaluate task performance as a function of the base model used. We employ two different inference techniques to generate the results reported here, namely using System-1 (direct answer) and System-2 (employing reasoning) thinking, as defined by [McGlynn2014ThinkingFA]. We do this by appending the tags "<answer>" or "<reasoning>" respectively upon generation, which induces models into either type of thinking. The results shown in Table 2 show the performances of our models across the multiple tasks defined in Section A, along with the diagnostic metrics defined in Section 3.1.
@@
Previous works have built models for similar tasks [zhang2024chemllmchemicallargelanguage, ether0, naturelm]. In Table 2, we compare against a baseline of similar size, ChemLLM, a model trained starting from InternLM2-Base-7B [internlm2]. Our results show that, on the tasks evaluated, ChemLLM is outperformed by our RL-tuned 3B models. The advantage is especially striking in the tasks of reaction naming and and reaction prediction. This contrasts with the results presented in the original publication, reporting over 88% 5-shot accuracy on retrosynthesis which could not be reproduced.
@@ 5.2 SCS as a predictive framework
A main claim of this paper is that it is important to develop diagnostic metrics to help prognosticate the performance of models before RL is applied, similar to the role of scaling laws in pre-training [kaplan2020scaling]. Here we show that pre-RL SCS is effective for this. As shown in Figure 4, pre-RL SCS reliably predicts post-RL success of models. It is observed that, on the tasks evaluated, low pre-RL SCS predictably yields incapable models, while high SCS leads to better models, especially when the models are trained with RL on the tasks being evaluated. We find strong correlations between pre-RL SCS and post-RL performance, namely $\rho=0.64$ for reaction prediction and $\rho=0.60$ for IUPAC translation across both 3B and 7B models. This provides a quantitative empirical framework to support future model development: SCS > 1.5 is a necessary threshold for RL to unlock reasoning.
@@
Qwen-2.5-14B [qwen25] 1.161
Qwen-2.5-32B [qwen25] 1.238
Qwen-2.5-72B [qwen25] 1.094
Llama3-70B [llama3] 1.481
Gemma-2-27B [gemma2] 2.097
Mistral-24B [mistral24] 2.324
@@
As case-study, we use SCS to retrospectively select the best base model to train ether0, the first chemical reasoning model [ether0]. We selected a range of open-source base LLMs across a variety of providers and sizes between 14B and 70B, and measured SCS. The results in Table 5 show that Mistral-24B would be the best option, followed by Gemma-2-27B, while the rest of models, even some beyond 70B in size, fall below the 1.5 SCS limit established previously. Within a similar size range, and despite its much better scores at most public benchmarks [mistral24], Qwen-2.5-32B is predicted to be a much worse base model for RL in chemistry tasks. These results match the base model selection from ether0 [ether0], that effectively used Mistral-24B as a base, further increasing confidence on the predictive power of measures of this type.